OpenStudy (anonymous):

Find the angles in the third and fourth quadrants to the nearest minute: sin(theta) = -0.0267

OpenStudy (jdoe0001):

\(\bf sin(\theta) = -0.0267\implies sin^{-1}[sin(\theta)]=sin^{-1}(-0.0267)\\ \quad \\ \theta=sin^{-1}(-0.0267)\)

OpenStudy (anonymous):

Thanks, but the answer it's giving me is -1.52998, and I don't know how I'm supposed to find it in the third and fourth quadrants.